OpenStudy (anonymous):

florida, california, and louisiana have hosted 35 super bowl games total. florida hosted 6 more games than louisiana. together, florida ansd louisiaana hosted 2 more than twice than california hosted. how many games did each state host?

OpenStudy (anonymous):

California = 11 Louisiana = 9 Florida = 15

OpenStudy (anonymous):

wow tht was fast..but how did u set yur equations up....i have to identify 3 variables and 3 equations....variables are easy but the equations are a lil harder...........

OpenStudy (anonymous):

actually that's a system of equations, just like what you are saying/

OpenStudy (anonymous):

California = x Louisana = y Florida = 35-x-y florida ansd louisiaana hosted 2 more than twice than california hosted so 35-x-y +y = 2y +2 35-x=2x+2 x=11 florida hosted 6 more games than louisiana so 35-x-y = y +6 putting x =11 y= 9 @ebydaqueen17 do the rest

OpenStudy (anonymous):

okay thanks